# Thermodynamic geometry in hadron resonance gas model at real and imaginary baryon chemical potential and a simple sufficient condition for quark deconfinement

### v2

@@ VII Summary
(1) In the case of the HRG model without EVE: The scalar curvature $R$ at real $\mu$ is negative only when $\mu$ and $T$ is small. This is consistent with the previous results Castorina et al. (2018, 2019). The curve of $R=0$ in the $\mu^{2}$-$T$ plane continues analytically from the imaginary chemical potential region to the real one.
(2) In the case of the HRG model with EVE: The scalar curvature $R$ at real $\mu$ is negative when $\mu$ and $T$ is small. This is consistent with the previous result Castorina et al. (2019). There are another regions of the negative $R$ when $T$ or $\mu$ is large. At $\mu=0$, $R$ vanishes when $T=T_{1}=T_{\rm RWL}=0.2103$ GeV and $T=T_{2}=0.161$ GeV. As is in the previous result Castorina et al. (2019), $T_{2}$ is consistent with the LQCD predicted pseudocritical temperature. In the wide region of imaginary $\mu$, $R$ is negative. However, at $\theta=\pi$, $R$ vanishes in the limit $T\to T_{\rm RWL}$ and there is another curve of $R=0$ in the vicinity of the RWL point. The sign changing of $R$ is related to the singularity at the RWL point. One of the curves of $R=0$ in the $\mu^{2}$-$T$ plane continues analytically from the imaginary chemical potential region to the real one. As is in the previous result Castorina et al. (2019), the part of this curve is consistent with the LQCD crossover line when $\mu<0.1$ GeV, but deviates from it when $\mu$ increases. On the other hand, the LQCD predicted critical point lies just below the lower curve of $R=0$ at $\mu\sim 0.6$ GeV.
(3) In the wide range of the imaginary chemical potential region, the determinant $g^{\prime}=-g$ of the metric of the $(\theta,T)$ coordinate is negative, hence, the system is expected to be thermodynamic unstable there. Nevertheless, the analytical continuation of the scalar curvature $R$ is possible from the imaginary $\mu$ region to the real one.
@@
We have also investigated the limitation temperature of the baryon gas model with EVE. The main results obtained here are summarized as follows.
@@
(8) The lower $T_{\rm max}(\mu)$ curve in the summary (6) is just below the curve of $n_{\rm b}=1/(2v_{\rm B})$. This fact means that $n_{\rm b}(\mu,T_{\rm max}(\mu))<1/(2v_{\rm B})$. When $\mu$ is large, $n_{\rm B}\sim n_{\rm b}$. Hence, we obtain a simple sufficient condition $n_{\rm B}>1/(2v_{\rm B})$ for quark deconfinement at large $\mu$, if we regard the lower $T_{\rm max}(\mu)$ as the limitation temperature of baryons.
In the HRG model with EVE, the curve of $R=0$ on the $\mu$-$T$ plane is consistent with the LQCD crossover line when $\mu<0.1$ GeV. However, the curve deviates from the LQCD crossover line as $\mu$ increases. On the other hand the LQCD crossover line meets the curve of $T_{\max,2}(\mu)$ at $\mu=0.3$ GeV. This indicates that the nature of the LQCD crossover line changes as $\mu$ increases. In fact, it was pointed out that, near the critical point, the ordering density is a linear combination of the scalar density, the net baryon number density and the energy density rather than the pure scalar density itself Fujii and Ohtani (2004). The effects of the baryon number density and the energy density may be more important than the scalar density itself in the large $\mu$ region. This may be the reason why the HRG model can predict the CP location approximately although the model has no mechanism of the chiral phase transition. Further studies to investigate the nature of the transition at CP is needed.

### v3

@@ VIII Summary
(1) In the case of the HRG model without EVE: The scalar curvature $R$ at real $\mu$ is negative only when $\mu$ and $T$ are small. This is consistent with the previous results Castorina et al. (2018, 2019). The curve of $R=0$ in the $\mu^{2}$-$T$ plane continues analytically from the imaginary chemical potential region to the real one.
(2) In the case of the HRG model with EVE: The scalar curvature $R$ at real $\mu$ is negative when $\mu$ and $T$ are small. This is consistent with the previous result Castorina et al. (2019). There are other regions of the negative $R$ when $T$ or $\mu$ is large. At $\mu=0$, $R$ vanishes when $T=T_{1}=T_{\rm RWL}=0.2103$ GeV and $T=T_{2}=0.161$ GeV. As is in the previous result Castorina et al. (2019), $T_{2}$ is consistent with the LQCD predicted pseudocritical temperature. In the wide region of imaginary $\mu$, $R$ is negative. However, at $\theta=\pi$, $R$ vanishes in the limit $T\to T_{\rm RWL}$ and there is another curve of $R=0$ in the vicinity of the RWL point. The sign change of $R$ is related to the singularity at the RWL point. One of the curves of $R=0$ in the $\mu^{2}$-$T$ plane continues analytically from the imaginary chemical potential region to the real one. As is in the previous study Castorina et al. (2019), the part of this curve is consistent with the LQCD crossover line when $\mu<0.1$ GeV, but deviates from it when $\mu$ increases. On the other hand, the LQCD predicted critical point lies just below the lower curve of $R=0$ at $\mu\sim 0.6$ GeV.
(3) In the wide range of the imaginary chemical potential region, the determinant $g^{\prime}=-g$ of the metric of the $(\theta,T)$ coordinate is negative, hence, the system is expected to be thermodynamically unstable there. Nevertheless, the analytical continuation of the scalar curvature $R$ is possible from the imaginary $\mu$ region to the real one.
@@
We have also investigated the limiting temperature of the baryon gas model with EVE. The main results obtained here are summarized as follows.
@@
(8) The lower $T_{\rm max}(\mu)$ curve in the summary (6) is just below the curve of $n_{\rm b}=1/(2v_{\rm B})$. This fact means that the limiting density $n_{\rm b}(\mu,T_{\rm max}(\mu))$ is smaller than $1/(2v_{\rm B})$. When $\mu$ is large, $n_{\rm B}\sim n_{\rm b}$. Hence, we obtain a simple empiric sufficient condition $n_{\rm B}>1/(2v_{\rm B})$ for quark deconfinement at large $\mu$, if we regard the global $T_{\rm max}(\mu)$ as the limiting temperature of baryons.
In the HRG model with EVE, the curve of $R=0$ on the $\mu$-$T$ plane is consistent with the LQCD crossover line when $\mu<0.1$ GeV. However, the curve deviates from the LQCD crossover line as $\mu$ increases. On the other hand the LQCD crossover line meets the curve of $T_{\max,2}(\mu)$ at $\mu=0.3$ GeV. This indicates that the nature of the LQCD crossover line changes as $\mu$ increases. In fact, it was pointed out that, near the critical point, the ordering density is a linear combination of the scalar density, the net baryon number density and the energy density rather than the pure scalar density itself Fujii and Ohtani (2004). The effects of the baryon number density and the energy density may be more important than the scalar density itself in the large $\mu$ region. This may be the reason why the HRG model can predict the CP location approximately although the model has no mechanism of the chiral phase transition. Further studies to investigate the nature of the transition at CP are needed.
@@
In this paper, we set $r_{\rm B}=0.8$ fm except for Sec. VII in which the $r_{\rm B}$-dependence of the results is investigated. It is shown that the qualitative features are robust under the moderate variation of $r_{\rm B}$. However, the quantitative features are somewhat changed. The baryon radius $r_{\rm B}$ is related to the Roberge-Weiss like temperature $T_{\rm RWL}$ which is related to the Roberge-Weiss temperature $T_{\rm RW}$. To eliminate the ambiguity of $r_{\rm B}$, more accurate determination of $T_{\rm RW}$ is needed.